OpenStudy (anonymous):

what is tanx=2cscx in the interval of -pi

OpenStudy (anonymous):

can u explain ur question in detail...are u asking for the value of tanx or x??

OpenStudy (anonymous):

here we can write tan x =sin x/cos x and csc x= 1/sin x then our equation will be sin x/cos x = 2/sin x =>(sin x)^2 =2cos x =>1-(cos x)^2 =2cos x {as u know (sin x)^2 + (cos x)^2 =1} =>(cos x)^2+ 2cos x-1=0 => t^2 +2t -1 =0 (let cos x=t) now we can find the soln of quadratic eqn t= {-2+ sqrt(4+4) }/2 or {-2- sqrt(4+4) }/2 = 0.414 or -2.414 but as we know the value of cos x cant be less than -1 ,the 2nd value(-2.414) is discarded... so we only have t= 0.414 => cos x = 0.414 => x = arc cos (0.414) => x = 1.144 so ans is d.